OpenStudy (anonymous):

What are the prime factors of 18? 2, 9 3, 6 2, 3, 6, 9 2, 3, 3

OpenStudy (anonymous):

2*3*3

OpenStudy (anonymous):

Which is the smallest number? zero point five two two-fifths eleven twenty-fifths nine-twentieths @gold_grace123

OpenStudy (anonymous):

two-fifths

OpenStudy (anonymous):

What is the greatest common factor of 16 and 48? 4 16 48 768 @gold_grace123

OpenStudy (anonymous):

16 is a factor of 48. Since the largest possible factor of 16 is 16 itself, the greatest common factor of 16 and 48 is 16. The factors of 16 are 1, 2, 4, 8, and 16. The factors of 48 are 1, 2, 3, 4, 6, 8, 12, 16, 24, and 48. The factors they have in common are 1, 2, 4, 8, and 16. Therefore, 16 is the greatest common factor. The greatest common factor can also be determined from the prime factors. The prime factors of 16 are 2, 2, 2, and 2. The prime factors of 48 are 2, 2, 2, 2, and 3. The prime factors they have in common are 2, 2, 2, and 2, so the greatest common factor is 2 x 2 x 2 x 2 = 16.

OpenStudy (anonymous):

To use mental math and compare five-sixths to one-half, which fraction would you use? five-halves three-sixths five-halves five-thirds @gold_grace123

OpenStudy (cj49):

to make 1/2=5/6 1/2x 5/3=5/6

OpenStudy (anonymous):

@sjman1 you put five-halves twice?

OpenStudy (anonymous):

Which shows forty-eight sixtieths in simplest form? four-thirds four-fifteenths eight-tenths four-fifths @gold_grace123

OpenStudy (cj49):

u mean 48/60

OpenStudy (anonymous):

12/15

OpenStudy (anonymous):

which equals 4/5

OpenStudy (anonymous):

Which of these is equivalent to seven-twelfths? twelve-nineteenths twenty-one fourths fourteen-nineteenths twenty-eight forty-eighths @gold_grace123

OpenStudy (anonymous):

twenty-eight forty-eighths

OpenStudy (anonymous):

Anything else? lol

OpenStudy (anonymous):

Which of the following is true? 0 is prime number 6 is a prime number twelve is a composite number twenty-three is a composite numbe @gold_grace123